23 June 2010

NEEM camp in full operation

The sun never sets on the sleeping tents during drilling season
NEEM city centre and suburbs. When camp population increases it becomes popular to move ‘out of town' in smaller tents.

After yesterdays crew exchange the camp population is now 37 persons which is about the upper limit for an operational camp. Everything is up and running: drilling progresses well towards the exciting Eemian ice which may be reached about 100 m below present drilling depth. Ice core processing is generally following the drilling, but at the moment there is time to process some of the Holocene brittle zone that was left in camp from last year to relax.
CFA (the continuous chemistry analysis) and all of the associated analyses have been making very good progress all season and right now 40,000 year old ice from the last glacial period is analyzed.

The skiway is being prepared and cargo is prepared for a flight operation tomorrow. Many camp activities are being filmed and documented by our outreach participants. Last but not least there is full activity in the kitchen all day long to prepare food for all the hard working people.

What we have done today:
1. Drilling and core logging.
2. Processed ice cores: 28 bags of brittle ice: bag 1117 to 1145.
3. CFA measurements: 8.8 m. Final depth: 1778.15 m.
4. Groomed skiway to prepare for Skier tomorrow.
5. Repaired one out of two broken spinners for ice chips.

Ad.1: Driller's report:
Over the last two days we continued to drill with the new cutters. They perform well with a pitch of 2.7 mm. Chips recovery is still less than optimal which sometimes leads to short runs after long runs. This can occur via ice collecting under the shoes or excess packing around the drill head early in the run.
On one of the hallow shafts we have again tried a configuration with a closed shaft and no lower valve at the super banger. At the moment it seems to work at least as well as the open shaft configuration and we hope it could provide some vacuum cleaning of extra chips if we suspect a dirty hole from a previous run.
The new teams are set with Sigfus, Trevor and Thorbjorn in the mornings and Romain, HC and Ji Woong in the evenings. And we have Sebastian in the wings if we need him.
We have had trouble with both centrifuge spinners over the last two days. Both were down for most of Tuesday and half of today. One is operational again and the other has seen its last days.

Driller's depth: 2193.11 m. Loggers depth: 2211.48 meters.

Weather: Blue sky and good visibility all day long. Temp. -13 °C to -6 °C, wind 3-10 knots from E.

FL, Anders Svensson


← Previous entry   Next entry →